OpenStudy (anonymous):

what is the absoulute value of |x-4|=|3-x|

OpenStudy (anonymous):

(x-4)^2 = (3-x)^2 -8x +16 = 9 - 6x -2x = -7 -> x= 7/2

OpenStudy (anonymous):

7/2